# Thermodynamic and energetic constraints on out-of-equilibrium tunneling rates

### v1

@@ V Conclusions
We have studied the tunneling rates of transitions between two weakly coupled, but otherwise generic, subsystems driven out of equilibrium by a temperature bias. For such setups, we have proved two novel bounds that the resulting out-of-equilibrium tunneling rates (6) satisfy: i) The thermodynamic constraint (16), stating that the out-of-equilibrium response of the tunneling rates bounds from below the magnitude of the rates themselves multiplied by the heat dissipated while establishing or depleting the temperature bias. ii) The energetic constraint (20), which states that the out-of-equilibrium response of the tunneling rates also bounds from below the tunneling rates weighted by the continuous change in the internal energy required to heat one of the subsystems. As a key consequence of these bounds, also the tunneling current and its low-frequency noise become bounded in the presence of a temperature bias.
@@
Our results thus highlight a fundamental connection between thermodynamic potentials and transport quantities for small-size quantum systems. In particular, they should be testable for a broad range of state-of-the-art experimental setups, including optical or mechanical cavities coupled together [61] or to (artificial) atoms [62, 63, 64, 65, 66, 67], and tunneling bridges across molecules [68, 38, 40] or magnetic impurities [69, 70, 71].
@@ Acknowledgements.
We gratefully acknowledge funding from the Knut and Alice Wallenberg foundation via the fellowship program (L.T. and J.S.), the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (101088169/NanoRecycle) (J.S.), from the Swedish Vetenskapsrådet via Project No. 2023-04043 (C.S), and from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101031655 (TEAPOT) (C.S.).
Appendix A Perturbative expansion of current and noise
In this Appendix, we re-derive the expressions (8) for the current $I$ and the noise $S$ obtained in Refs. [17, 24, 26]. As our starting point, we consider the unitary evolution from time 00 to time $t$ of the full Hamiltonian $\hat{H}(t)=\hat{H}_{0}+\hat{V}(t)$. The time evolution operator reads
[…]
$\hat{\mathcal{U}}_{0}(t,0)$ $\equiv e^{-i\hat{H}_{0}t/\hbar},$ (29a)
[…]
are the evolution induced by the free Hamiltonian $\hat{H}_{0}$ and the first correction due to the tunneling Hamiltonian $\hat{V}(t)$, respectively.
With Eq. (29), also the average current can be expanded in powers of the tunneling operator $A$ as $\langle\hat{I}_{\text{H}}(t)\rangle\approx I^{(1)}(t)+I^{(2)}(t)$ with
$I^{(2)}(t)$ $\equiv\text{Tr}\left\{[\delta\mathcal{U}^{\dagger}(t,0)\hat{I}(t)\hat{\mathcal{U}}_{0}(t,0)+\hat{\mathcal{U}}_{0}^{\dagger}\hat{I}(t)\delta\mathcal{U}(t,0)]\hat{\rho}_{0}\right\},$ (30b)
with $\hat{I}(t)$ given in Eq. (7). In Eq. (30), $\hat{\rho}_{0}$ is the state of the system at time $t=0$. In the following, we assume that $[\hat{\rho}_{0},\hat{H}_{0}]=0$, such that, if {|$\{\ket{a}\}$ are the eigenstates of $\hat{H}_{0}$ with energies $\epsilon_{a}$, their occupations are given by $\hat{\rho}_{0}\ket{a}=p_{a}\ket{a}$.
By defining the matrix elements $A_{ab}\equiv\braket{a}{\hat{A}}{b}$, $A^{\dagger}_{ab}\equiv\braket{a}{\hat{A}^{\dagger}}{b}=A_{ba}^{*}$ and the energy difference $\epsilon_{ba}\equiv\epsilon_{b}-\epsilon_{a}$, the first order contribution to the current (30a) becomes
$I^{(1)}(t)=-\frac{iq}{\hbar}\sum_{a}\left(e^{-i\omega t}A_{aa}-e^{i\omega t}A^{\dagger}_{aa}\right)p_{a},$ (31)
and the second order contribution (30b) becomes
$\begin{split}I^{(2)}(t)&=\frac{q}{\hbar^{2}}\int_{0}^{t}dt^{\prime}\sum_{ab}p_{a}\left\{\left[A_{ab}A_{ba}e^{-i\omega(t^{\prime}+t)}-\text{c.c.}\right]\times\right.\\ &\times 2i\sin[\frac{\epsilon_{ba}}{\hbar}(t-t^{\prime})]+|A_{ba}|^{2}2\cos[(\frac{\epsilon_{ba}}{\hbar}-\omega)(t-t^{\prime})]\\ &\left.-|A_{ab}|^{2}2\cos[(\frac{\epsilon_{ba}}{\hbar}+\omega)(t-t^{\prime})]\right\}.\end{split}$ (32)
Now, since by hypothesis we have $[\hat{Q},\hat{A}]=q\hat{A}$ and $[\hat{Q},\hat{H}_{0}]=0$, we find that
$\braket{a}{[\hat{Q},\hat{A}]}{b}=(q_{a}-q_{b})A_{ab}=qA_{ab},$ (33)
where $\hat{Q}\ket{a}=q_{a}\ket{a}$. This identity implies that $A_{aa}=0$ (as long as $q\neq 0$), and that $A_{ab}A_{ba}=0$. Thus, the first order contribution to the current (31) vanishes, $I^{(1)}(t)=0$, and the second order contribution (32) simplifies to
$\begin{split}&I^{(2)}(t)=\frac{q}{\hbar^{2}}\sum_{ab}p_{a}\left\{|A_{ba}|^{2}\frac{2\sin[(\frac{\epsilon_{ba}}{\hbar}-\omega)t]}{\frac{\epsilon_{ba}}{\hbar}-\omega}\right.\\ &\qquad\qquad\qquad\qquad\left.-|A_{ab}|^{2}\frac{2\sin[(\frac{\epsilon_{ba}}{\hbar}+\omega)t]}{\frac{\epsilon_{ba}}{\hbar}+\omega}\right\},\\ &\stackrel{{\scriptstyle t\to\infty}}{{\to}}\frac{2\pi q}{\hbar^{2}}\sum_{ab}p_{a}\left\{|A_{ba}|^{2}\delta(\frac{\epsilon_{ba}}{\hbar}-\omega)-|A_{ab}|^{2}\delta(\frac{\epsilon_{ba}}{\hbar}+\omega)\right\},\end{split}$ (34)
which is identical to Eq. (8a).
To calculate the noise we start from the two-time correlator $\mathcal{S}(t)\equiv\langle\hat{I}_{\text{H}}(t)\hat{I}(0)\rangle-\langle\hat{I}_{\text{H}}(t)\rangle\langle\hat{I}(0)\rangle$, which at lowest order in the tunneling operator $\hat{A}$ reads
$\begin{split}\mathcal{S}(t)&\approx\langle\hat{I}_{\text{H}}(t)\hat{I}(0)\rangle\approx\text{Tr}\left\{\hat{\mathcal{U}}_{0}^{\dagger}(t,0)\hat{I}(t)\hat{\mathcal{U}}_{0}(t,0)\hat{I}(0)\hat{\rho}_{0}\right\}\\ &\approx-\frac{q^{2}}{\hbar^{2}}\sum_{ab}e^{-i\epsilon_{ba}t/\hbar}p_{a}\left(e^{-i\omega t}A_{ab}-e^{i\omega t}A_{ab}^{\dagger}\right)\times\\ &\hskip 113.81102pt\times\left(A_{ba}-A_{ba}^{\dagger}\right).\end{split}$ (35)
[…]
$\mathcal{S}\approx\frac{2\pi q}{\hbar^{2}}\sum_{ab}p_{a}\left\{|A_{ba}|^{2}\delta(\frac{\epsilon_{ba}}{\hbar}-\omega)+|A_{ab}|^{2}\delta(\frac{\epsilon_{ba}}{\hbar}+\omega)\right\}.$ (36)
This expression is equivalent to Eq. (8b).
We finally note that, by exchanging index labels, the current and the noise can we re-written as
$I$ $\approx\frac{2\pi q}{\hbar^{2}}\sum_{ab}|A_{ba}|^{2}\delta(\frac{\epsilon_{ba}}{\hbar}-\omega)\left[p_{a}-p_{b}\right],$ (37a)
$S$ $\approx\frac{2\pi q^{2}}{\hbar^{2}}\sum_{ab}|A_{ba}|^{2}\delta(\frac{\epsilon_{ba}}{\hbar}-\omega)\left[p_{a}+p_{b}\right].$ (37b)
If the system is initially, at time 0, in equilibrium at the inverse temperature $\beta$, the initial state is a Gibbs state $\hat{\rho}_{0}\propto e^{-\beta\hat{H}_{0}}$ and the state occupation probabilities are given as $p_{a}\propto e^{-\beta\epsilon_{a}}$. In this case, the products in the summands in Eq. (37) satisfy
$\delta(\frac{\epsilon_{ba}}{\hbar}-\omega)[p_{a}\pm p_{b}]=\delta(\frac{\epsilon_{ba}}{\hbar}-\omega)\left[e^{-\beta\hbar\omega}\pm 1\right]p_{b},$ (38)
establishing the equal temperature relation between current and noise, given in Eq. (10).
Appendix B More general tunneling operators
When more generic tunneling operators are considered, the bounds derived in the main text remain valid for the individual tunneling rates $\Gamma_{\rightleftarrows}$. They do however not extend to the tunneling current and the tunneling current noise.
To show this, in the following, we thus drop the condition $[\hat{Q},\hat{A}]=q\hat{A}$ on the tunneling operator $\hat{A}$. Even without this condition, it is possible to define a current operator via Heisenberg’s equation of motion as
$\hat{I}(t)\equiv-\frac{i}{\hbar}\left([\hat{Q},\hat{A}]e^{-i\omega t}+[\hat{Q},\hat{A}^{\dagger}]e^{i\omega t}\right).$ (39)
Analogously to the derivation in Appendix A, we calculate the lowest-order contributions in the tunneling operator $\hat{A}$, which now does not satisfy any commutation property with the observable $\hat{Q}$, but we still assume that it satisfies $A_{ab}A_{ba}=0$. This constraint means that only one of the transitions between |$\ket{a}\to\ket{b}$ and |$\ket{b}\to\ket{a}$ is possible, where |$\ket{a}$ and |$\ket{b}$ are the common eigenstates of $\hat{H}_{0}$ and $\hat{Q}$ (recall that $[\hat{H_{0}},\hat{Q}]=0$). The forward and the backward rates between these eigenstates are now given as
$\begin{split}\Gamma_{ba}^{\rightarrow}&\equiv\frac{2\pi}{\hbar}\delta(\epsilon_{ba}-\hbar\omega)|A_{ba}|^{2}p_{a},\\ \Gamma_{ba}^{\leftarrow}&\equiv\frac{2\pi}{\hbar}\delta(\epsilon_{ba}-\hbar\omega)|A_{ba}|^{2}p_{b}.\\ \end{split}$ (40)
With these rates, the current and the noise read
$\begin{split}I&=\sum_{ab}(q_{b}-q_{a})[\Gamma^{\rightarrow}_{ba}-\Gamma^{\leftarrow}_{ba}],\\ \mathcal{S}&=\sum_{ab}(q_{b}-q_{a})^{2}[\Gamma^{\rightarrow}_{ba}+\Gamma^{\leftarrow}_{ba}],\\ \end{split}$ (41)
where we recall that $q_{a}\ket{a}=\hat{Q}\ket{a}$. Specializing to a bipartite system, the rates take the form
$\begin{split}\Gamma_{mk,nl}^{\rightarrow}&\equiv\frac{2\pi}{\hbar}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega)|A_{mk,nl}|^{2}p_{n}^{\text{(L)}}p_{l}^{\text{(R)}},\\ \Gamma_{mk,nl}^{\leftarrow}&\equiv\frac{2\pi}{\hbar}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega)|A_{mk,nl}|^{2}p_{m}^{\text{(L)}}p_{k}^{\text{(R)}}.\\ \end{split}$ (42)
By inspecting the proofs in Appendix D and F, we find that both the thermodynamic and energetic constraints presented in the main text remain valid for the rates $\Gamma_{mk,nl}^{\rightleftarrows}$ between each pair of distinct eigenstates. If the hypothesis $[\hat{Q},\hat{A}]=q\hat{A}$ is restored, one has $(q_{mk}-q_{nl})\to q$, and the results in the main text are thus recovered.
Using these rates, the current and noise become
$\begin{split}I&=\sum_{mnlk}(q_{mk}-q_{nl})[\Gamma^{\rightarrow}_{mk,nl}-\Gamma^{\leftarrow}_{mk,nl}],\\ \mathcal{S}&=\sum_{mnlk}(q_{mk}-q_{nl})^{2}[\Gamma^{\rightarrow}_{mk,nl}+\Gamma^{\leftarrow}_{mk,nl}].\\ \end{split}$ (43)
Here, $q_{nl}\ket{nl}=\hat{Q}\ket{nl}$, and the eigenstates |$\ket{nl}$ satisfy $\hat{H}_{\text{L}}\ket{nl}=\epsilon_{n}^{(\text{L})}\ket{nl}$, $\hat{H}_{\text{R}}\ket{nl}=\epsilon_{l}^{(\text{R})}\ket{nl}$. Therefore, we see that the only difference compared to Eq. (8) in the main text is that the rates at different resonances are weighted with state-dependent factors $(q_{mk}-q_{nl})$ and $(q_{mk}-q_{nl})^{2}$, instead of the simpler factors $q$ and $q^{2}$. This feature prevents us from defining common global rates with which both the current and the noise can be expressed.
Appendix C Work extraction and nonequilibrium free energy
@@
$=-\text{Tr}\left\{\left(\hat{H}_{\text{S}}(t)+\hat{H}_{\text{SB}}(t)\right)\frac{d\hat{\rho}(t)}{dt}\right\},$
respectively. In this way, we can write an expression equivalent to the first law of thermodynamics. To this end, one considers the change in the energy of S, denoted $U(t)\equiv\text{Tr}\left\{\left(\hat{H}_{\text{S}}(t)+\hat{H}_{\text{SB}}(t)\right)\hat{\rho}(t)\right\}$, namely
@@
where $T_{\text{B}}$ is the temperature of the bath. Then, combining Eqs. (47, 49, 50) we find that the performed work is limited from below by the nonequilibrium free energy $\Delta F$ as
@@
In this work, we are interested in the situation where, at the beginning and at the end of the operation, the system Hamiltonian and the system-bath interaction satisfy
@@
i.e. the system’s Hamiltonian is left unchanged after the operation, and the system and the bath are not coupled at the beginning and at the end of the operation. Then, the nonequilibrium free energy reads
$\Delta F=\Delta F_{\text{S}}=\Delta U_{\text{S}}(\tau)-T_{\text{B}}\Delta S_{\text{S}}(\tau)$ (53)
@@
We next combine Eqs. (47, 49, 56) in two different ways, depending on whether we focus on the heat absorbed by the bath, $\Delta Q_{\text{B}}(\tau)$, or on the energy variation in the system $\Delta U_{\text{S}}(\tau)$. For these two situations, we find
@@
As detailed in Fig. 1, we are interested in both cases $T<T_{\text{B}}$ and $T>T_{\text{B}}$, depending on whether we are cooling or heating subsystem L. When $T<T_{\text{B}}$, we focus on Eq. (57a). There, $\Delta U_{\text{S}}(\tau)\geq 0$, since the system is heated by the hot bath, and $W(\tau)\leq 0$ since we are using the heat flow to extract energy. Then
@@
Appendix D Derivation of the thermodynamic constraint
In this Appendix, we prove Eq. (16) for the forward tunneling rate $\Gamma_{\rightarrow}$ in Eq. (6). The proof for $\Gamma_{\leftarrow}$ is analogous. The difference between out-of-equilibrium rates $\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})$ and $\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})$ reads
$\begin{split}&\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})-\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})=\frac{2\pi}{\hbar}\sum_{nmlk}|\mathcal{A}_{nmlk}|^{2}\times\\ &\times\delta(\epsilon^{\text{(L)}}_{mn}+\epsilon^{\text{(R)}}_{kl}-\hbar\omega)\left[\frac{e^{-\beta_{\text{L,1}}\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,1}})}-\frac{e^{-\beta_{\text{L,2}}\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,2}})}\right]\frac{e^{-\beta_{\text{R}}\epsilon_{l}^{\text{(R)}}}}{Z_{\text{R}}(\beta_{\text{R}})}.\end{split}$ (60)
Here, the difference between the occupation probabilities of subsystem L vanishes at the special energy
$\tilde{\epsilon}^{\text{(L)}}=\frac{1}{\beta_{\text{L,1}}-\beta_{\text{L,2}}}\log\frac{Z_{\text{L}}(\beta_{\text{L,2}})}{Z_{\text{L}}(\beta_{\text{L,1}})}.$ (61)
When $\beta_{\text{L,1}}\geq\beta_{\text{L,2}}$, we then find that the following inequality holds:
$\begin{split}&e^{\alpha\epsilon_{n}^{\text{(L)}}}\left(\frac{e^{-\beta_{\text{L,1}}\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,1}})}-\frac{e^{-\beta_{\text{L,2}}\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,2}})}\right)\leq\\ &\hskip 56.9055pt\leq e^{\alpha\tilde{\epsilon}^{\text{(L)}}}\left(\frac{e^{-\beta_{\text{L,1}}\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,1}})}-\frac{e^{-\beta_{\text{L,2}}\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,2}})}\right)\end{split}$ (62)
for any real number $\alpha\geq 0$ and for all $\epsilon_{n}^{\text{(L)}}$. Note that this relation becomes an equality for $\epsilon_{n}^{\text{(L)}}=\tilde{\epsilon}^{\text{(L)}}$. Saturating this inequality is necessary to obtain a nontrivial equality in the thermodynamic constraint. Specifically, such a condition is achieved with a tunneling operator $\hat{A}$ allowing transitions starting from states of subsystem L that have this specific energy $\epsilon_{n}^{\text{(L)}}=\tilde{\epsilon}^{\text{(L)}}$.
Using this inequality for all resonances in Eq. (60), we have
$\begin{split}&\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})-\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})\leq\frac{2\pi}{\hbar}e^{\alpha\tilde{\epsilon}^{\text{(L)}}}\times\\ &\qquad\times\sum_{nmlk}|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon^{\text{(L)}}_{mn}+\epsilon^{\text{(R)}}_{kl}-\hbar\omega)\times\\ &\times\left[\frac{e^{-(\beta_{\text{L,1}}+\alpha)\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,1}})}-\frac{e^{-(\beta_{\text{L,2}}+\alpha)\epsilon_{n}^{\text{(L)}}}}{Z_{\text{L}}(\beta_{\text{L,2}})}\right]\frac{e^{-\beta_{\text{R}}\epsilon_{l}^{\text{(R)}}}}{Z_{\text{R}}(\beta_{\text{R}})}.\end{split}$ (63)
Here, we recognize the occupation probabilities at different temperatures, which allows us to write
$\begin{split}&\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})-\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})\leq e^{\alpha\tilde{\epsilon}^{\text{(L)}}}\times\\ &\qquad\times\left[\frac{Z_{\text{L}}(\beta_{\text{L,1}}+\alpha)}{Z_{\text{L}}(\beta_{\text{L,1}})}\Gamma_{\rightarrow}(\beta_{\text{L,1}}+\alpha,\beta_{\text{R}})+\right.\\ &\left.\qquad\qquad-\frac{Z_{\text{L}}(\beta_{\text{L,2}}+\alpha)}{Z_{\text{L}}(\beta_{\text{L,2}})}\Gamma_{\rightarrow}(\beta_{\text{L,2}}+\alpha,\beta_{\text{R}})\right]\end{split}$ (64)
for all $\alpha\geq 0$ and $\beta_{\text{L,1}}>\beta_{\text{L,2}}$. To avoid dealing with too many different temperatures, we study how Eq. (64) behaves as $\alpha/\beta_{\text{L,$i$/R}}\to 0^{+}$. In particular, to zeroth order in $\alpha$, the right-hand side cancels with the left-hand side. The first-order contribution in $\alpha$ leads instead to the inequality
$\begin{split}&\tilde{\epsilon}^{\text{(L)}}\left[\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})-\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})\right]+\\ &\qquad+[\partial_{\beta}\log Z_{\text{L}}(\beta_{\text{L,1}})]\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})+\\ &\qquad-[\partial_{\beta}\log Z_{\text{L}}(\beta_{\text{L,1}})]\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})\geq\\ &\qquad\qquad\geq\partial_{\text{L}}\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})-\partial_{\text{L}}\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}}),\end{split}$ (65)
where $\partial_{\text{L}}\Gamma_{\rightarrow}(x,y)\equiv\frac{\partial\Gamma_{\rightarrow}}{\partial x}(x,y)$. Here, we already recognize the structure of the rate response on the right-hand side of the inequality. Recalling that $\partial_{\beta}\log Z_{\text{L}}(\beta)=-U_{\text{L}}(\beta)$, we see that the rates are multiplied by the differences
$\begin{split}\tilde{\epsilon}^{\text{(L)}}-U_{\text{L}}(\beta_{\text{L,1}})&=\frac{\Delta S_{\text{L}}-\beta_{\text{L,2}}\Delta U_{\text{L}}}{\beta_{\text{L,1}}-\beta_{\text{L,2}}}=-\frac{\beta_{\text{L,2}}\Delta F_{\text{L}}^{\text{(h)}}}{\beta_{\text{L,1}}-\beta_{\text{L,2}}},\\ \tilde{\epsilon}^{\text{(L)}}-U_{\text{L}}(\beta_{\text{L,2}})&=\frac{\Delta S_{\text{L}}-\beta_{\text{L,1}}\Delta U_{\text{L}}}{\beta_{\text{L,1}}-\beta_{\text{L,2}}}=\frac{\beta_{\text{L,1}}\Delta F_{\text{L}}^{\text{(c)}}}{\beta_{\text{L,1}}-\beta_{\text{L,2}}}\end{split}$ (66)
where $\Delta U_{\text{L}}\equiv U_{\text{L}}(\beta_{\text{L,2}})-U_{\text{L}}(\beta_{\text{L,1}})$ and $\Delta S_{\text{L}}\equiv S_{\text{L}}(\beta_{\text{L,2}})-S_{\text{L}}(\beta_{\text{L,1}})$. The nonequilibrium free energy variations $\Delta F_{\text{L}}^{\text{(h/c)}}$ are defined with respect to the hot (cold) temperature $\beta_{\text{L,2}}^{-1}$ ($\beta_{\text{L,1}}^{-1}$) as
$\begin{split}\Delta F_{\text{L}}^{\text{(h)}}&\equiv\Delta U_{\text{L}}-T_{\text{L,2}}\Delta S_{\text{L}},\\ \Delta F_{\text{L}}^{\text{(c)}}&\equiv-\Delta U_{\text{L}}+T_{\text{L,1}}\Delta S_{\text{L}}.\\ \end{split}$ (67)
Then, combining Eqs. (65) and (66), we find
$\begin{split}&-\frac{T_{\text{L,1}}\Delta F_{\text{L}}^{\text{(h)}}}{T_{\text{L,2}}-T_{\text{L,1}}}\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}})-\frac{T_{\text{L,2}}\Delta F_{\text{L}}^{\text{(c)}}}{T_{\text{L,2}}-T_{\text{L,1}}}\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})\geq\\ &\qquad\geq\partial_{\text{L}}\Gamma_{\rightarrow}(\beta_{\text{L,2}},\beta_{\text{R}})-\partial_{\text{L}}\Gamma_{\rightarrow}(\beta_{\text{L,1}},\beta_{\text{R}}),\end{split}$ (68)
which holds for $\beta_{\text{L,1}}>\beta_{\text{L,2}}$. By choosing $\beta_{\text{L,1}}=\beta_{\text{R}}$ and any hotter inverse temperature as $\beta_{\text{L,2}}=\beta_{\text{L}}$, we arrive at Eq. (16).
Notably, the derivatives in the rate response $\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows}$ [Eq. (17)] can alternatively be understood in terms of higher-order correlation functions of the tunneling operator $\hat{A}$ in (5). More specifically, we find that

### v2

@@ IV Conclusions
We have studied the transition probabilities of two-point measurement schemes for different initial thermal state occupations, in an otherwise generic quantum system. For such setups, we have proved two novel bounds that the transition probabilities (2) satisfy: i) The thermodynamic constraint (23a), stating that the response of the transition probabilities for different initial temperatures bounds from below the magnitude of the probabilities themselves multiplied by the heat dissipated while cooling or heating the system to these temperatures. ii) The energetic constraint (23b), which states that the response of the transition probabilities to temperature variations also bounds from below the transition probabilities weighted by the continuous change in the internal energy required to heat the system.
As one application of interest, we have analyzed the developed constraints for the tunneling rates of transitions between two weakly coupled subsystems driven out of equilibrium by a temperature bias. As a key consequence of these bounds, also the tunneling current and its low-frequency noise become bounded in the presence of a temperature bias.
@@
Our results thus highlight a fundamental connection between thermodynamic potentials and transport quantities for small-size quantum systems. In particular, they should be testable for a broad range of state-of-the-art experimental setups, including optical or mechanical cavities coupled together [46] or to (artificial) atoms [47, 48, 49, 50, 51, 52], and tunneling bridges across molecules [53, 37, 39] or magnetic impurities [54, 55, 56]. Beyond the weak-coupling examples studied in the present paper, also the strong coupling regime is captured by our general constraints and it would be intriguing to investigate, e.g., coherent oscillations between strongly coupled few-level systems in the future.
@@ Acknowledgements.
We gratefully acknowledge funding from the Knut and Alice Wallenberg foundation via the fellowship program (L.T. and J.S.), the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (101088169/NanoRecycle) (J.S.), from the PNRR MUR project No. PE0000023-NQSTI (M.A.), from the Swedish Vetenskapsrådet via Project No. 2023-04043 (C.S.), and from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101031655 (TEAPOT) (C.S.).
Appendix A Work extraction and nonequilibrium free energy
@@
$=-\text{Tr}\left\{\left(\hat{H}_{\text{S}}(t)+\hat{H}_{\text{SB}}(t)\right)\frac{d\hat{\rho}(t)}{dt}\right\}.$
In this way, we can write an expression equivalent to the first law of thermodynamics. To this end, one considers the change in the energy of S, denoted $U(t)\equiv\text{Tr}\left\{\left(\hat{H}_{\text{S}}(t)+\hat{H}_{\text{SB}}(t)\right)\hat{\rho}(t)\right\}$, namely
@@
where $T_{\text{B}}$ is the temperature of the bath. Then, combining Eqs. (36, 38, 39) we find that the performed work is limited from below by the nonequilibrium free energy $\Delta F$ as
@@
In the present paper, we are interested in the situation where, at the beginning and at the end of the operation, the system Hamiltonian and the system-bath interaction satisfy
@@
i.e., the system Hamiltonian is left unchanged after the operation, and no system-bath coupling exists at the beginning and at the end of the operation. Then, the nonequilibrium free energy reads
$\Delta F=\Delta F_{\text{S}}=\Delta U_{\text{S}}(\tau)-T_{\text{B}}\Delta S_{\text{S}}(\tau),$ (42)
@@
We next combine Eqs. (36, 38, 45) in two different ways, depending on whether we focus on the heat absorbed by the bath, $\Delta Q_{\text{B}}(\tau)$, or on the energy variation in the system $\Delta U_{\text{S}}(\tau)$. For these two situations, we find
@@
As detailed in Fig. 2, we are interested in both cases $T<T_{\text{B}}$ and $T>T_{\text{B}}$, depending on whether we are cooling or heating the system. When $T<T_{\text{B}}$, we focus on Eq. (46a). There, $\Delta U_{\text{S}}(\tau)\geq 0$, since the system is heated by the hot bath, and $W(\tau)\leq 0$ since we are using the heat flow to extract energy. Then
@@
Appendix B Sufficient condition for trivial constraints
Both thermodynamic and energetic costs, Eqs. (14) and (18) respectively, are positive by definition, and are lower-bounded by the response term $\mathcal{W}^{\text{(Resp)}}_{a\to b}$ in Eq. (13). Therefore, the thermodynamic and energetic constraints are nontrivial whenever $\mathcal{W}^{\text{(Resp)}}_{a\to b}>0$. In this Appendix we provide a sufficient condition for $\mathcal{W}^{\text{(Resp)}}_{a\to b}<0$, making the results of Eqs. (23a, 23b) trivial, and a sufficient condition for $\mathcal{W}^{\text{(Resp)}}_{a\to b}>0$, making the results of Eqs. (23a, 23b) nontrivial. Starting from the definition of the response contribution (13) and recalling that derivatives with respect to $\beta$ only act on the probabilities $p_{a}$, we consider
$\partial_{\beta}p_{a}(\beta)=\left[U(\beta)-\epsilon_{a}\right]p_{a}(\beta).$ (49)
Taking a second derivative with respect to $\beta$ we find
$\partial_{\beta}^{2}p_{a}(\beta)=\left[\partial_{\beta}U(\beta)+[U(\beta)-\epsilon_{a}]^{2}\right]p_{a}(\beta).$ (50)
We now focus on all energies $\epsilon_{a}$ lying far away from the internal energy $U$. Namely, we suppose that the distance between $\epsilon_{a}$ and $U$ is larger than the standard deviation of the internal energy, such that the following inequality holds for all $\beta\in[\beta_{\mathrm{h}},\beta_{\mathrm{c}}]$,
$|U(\beta)-\epsilon_{a}|\geq\delta U(\beta),$ (51)
where
$\delta U(\beta)=\sqrt{\sum_{a}(\epsilon_{a})^{2}p_{a}(\beta)-U^{2}(\beta)}=\sqrt{-\partial_{\beta}U(\beta)}$ (52)
is the standard deviation of the internal energy. In this energy interval, defined by Eq. (51), the second derivative of the occupation probabilities is positive, i.e. $\partial^{2}_{\beta}p_{a}(\beta)\geq 0$, making $\partial_{\beta}p_{a}(\beta)$ an increasing function in the inverse temperature $\beta$. This in turn makes the rate response negative,
$\partial_{\beta}p_{a}(\beta_{\text{h}})-\partial_{\beta}p_{a}(\beta_{\text{c}})\geq 0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{a\to b}\leq 0,$ (53)
thus trivializing both thermodynamic and energetic constraints.
Conversely, the condition that the energies of the initial states lie within a standard deviation from the internal energy for all inverse temperatures in $[\beta_{\text{h}},\beta_{\text{c}}]$ is a sufficient condition for the rate response to be positive $\mathcal{W}^{\text{(Resp)}}_{a\to b}\geq 0$. Indeed, this condition on the energies makes $\partial_{\beta}p_{a}(\beta)$ a decreasing function in the inverse temperature $\beta$, leading to
$\partial_{\beta}p_{a}(\beta_{\text{h}})-\partial_{\beta}p_{a}(\beta_{\text{c}})\leq 0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{a\to b}\geq 0.$ (54)
We demonstrate this reasoning with a simple example: for the two-level system considered in Sec. III.3, we can easily compute both the internal energy and its standard deviation
Appendix C Transition rates in the weak-tunneling regime
In this Appendix, we derive expressions for the weak-tunneling transition rates of Eq. (21) starting from the two-point measurement probabilities of Eq. (2), closely following the derivation presented in [62]. As our starting point, we consider the unitary evolution from time 00 to time $t$ of the full Hamiltonian $\hat{H}(t)=\hat{H}_{0}+\hat{V}(t)$. The time-evolution operator reads